# Event2Vec: Processing Neuromorphic Events directly by Representations in Vector Space

### v3

@@ 3.4 Event Sampling and Aggregation
In this paper, we primarily use two methods. The first is uniform random sampling. We find that this straightforward method works well in most cases and is extremely computationally efficient. However, a significant limitation of random sampling is the substantial information loss incurred by discarding the majority of the events, leading to suboptimal accuracy in complex tasks. Our second method addresses this by leveraging k-means clustering to aggregate the entire event stream into $L$ representative clusters. Specifically, the clustering process is performed independently on the two event polarities to preserve their distinct information channels. After clustering, the centroid of each cluster is treated as a representative event. Its timestamp is taken directly from the centroid’s temporal coordinate, while its spatial coordinates are quantized to the nearest integers. Furthermore, we compute an intensity factor, $\rho$, equal to the number of raw events belonging to that cluster. This intensity factor then modulates the corresponding spatial embedding vector, effectively weighting the representation by its event density. This approach ensures that information from every event contributes to the final representation.
@@
$\textbf{V}[i]=\bm{\rho}[i]\cdot\text{Embed}_{s}(\textbf{x}[i],\textbf{y}[i],\textbf{p}[i])+\text{Embed}_{t}(\Delta\textbf{t})[i],$ (7)
where $\bm{\rho}$, x, y, p, and t are vectors representing the intensity factors, spatial coordinates, and timestamps for the sequence of $L$ events. The vector of temporal differences is defined as $\Delta\textbf{t}[i]=\textbf{t}[i+1]-\textbf{t}[i]$, with the final value set to zero. For a native event, $\bm{\rho}[i]$ is 1, while for a cluster event, it represents the number of raw events aggregated into that cluster.
@@ 3.5 Network structure
In this paper, we employ the Transformer architecture, leveraging its core strengths: the ability to efficiently process sequences in parallel and capture long-range dependencies. These characteristics make it particularly well-suited for the sequential representations generated by event2vec.
As shown in Figure 2(a), the spatial embedding module $\phi$ consists of a stack of linear layers. It gradually increases the number of features from 3 to $D$. Layer Normalization (Ba et al., 2016) layers are also inserted to stabilize training but omitted from the figure for clarity. The temporal embedding module has a similar structure to the spatial embedding module, except that it uses depth-wise convolutional layers with a kernel size of 3 and a stride of 1, shown in Figure 2(b).
We employ the Forgetting Transformer (Lin et al., 2025) as the linear attention in Figure 2(c). It is important to recognize that linear attention is fundamentally equivalent to RNNs (Katharopoulos et al., 2020), operating with fixed-size hidden states. To enhance the learning capability for extremely long event sequences, we extend the linear attention model to a parameter-shared bi-directional formulation. Further details are provided in Appendix A.2. A linear attention module and a feedforward network, consisting of two linear layers, together constitute a single backbone block. As illustrated in Figure 2(c), the full backbone is composed of $n$ such blocks stacked sequentially. For the classification head, we employ an average pooling layer to aggregate features across all positions in the sequence.
@@ 4 Experiments
To validate the event2vec representation, we conduct a series of experiments on classification tasks using three neuromorphic datasets: DVS Gesture (Amir et al., 2017), ASL-DVS (Bi et al., 2019), and DVS-Lip (Tan et al., 2022). In this section, results are reported in the format $a\pm b$, representing the mean and standard deviation, respectively. For experiments that involve random sampling, these statistics are computed over 10 independent runs on the test set. Experimental details are provided in Appendices A.3 to A.8.
@@ 4.1 Comparison between representations
Accuracy and Parameter Efficiency Table 1 compares the accuracy and model parameters of event2vec with those of other representations across the three datasets. Our models for DVS Gesture and ASL-DVS are trained directly on randomly sampled events. For DVS-Lip, our model first undergoes self-supervised pre-training (refer to Appendix A.5) on cluster events. We then report the fine-tuning accuracy on both randomly sampled events and cluster events. Our method achieves accuracy comparable to that of other leading representations while demonstrating exceptional parameter efficiency.
Table 1: Model performance and size comparison on neuromorphic datasets
@@
DVS Gesture Sparse GRU + Frame (Subramoney et al., 2023) 97.80 4.8
SNN + Frame (Yao et al., 2023) 98.23 6.5
@@
GNN,Transformer + Image,Voxel Graph (Yuan et al., 2023) 99.60 220.3
@@
DVS-Lip ResNet-18,BiGRU + Frame (Tan et al., 2022) 60.3 72.1
Spiking ResNet18,BiGRU + Frame (Dampfhoffer & Mesquida, 2024) 75.3 58.6
@@
Linear Attention + Event2vec (1024 Cluster Events) 75.14 18.3
Table 2: Comparative analysis of pre-processing latency for event representations on DVS Gesture
Representation Hyper-Parameter Total Pre-processing Time (s)
@@
[…]
Table 3: Benchmark of throughput and single event stream inference latency on DVS Gesture
Method Throughput (samples/s) Single Stream Latency (ms)
Training Inference
[…]
Event Pre-processing Time Table 2 benchmarks the total pre-processing time required by different representations to process all 1,176 event streams of the DVS Gesture training set. Hyper-parameters were set to the values specified in their respective original publications, where available. For others, we adopted commonly used values; for instance, graph construction was constrained to an average node degree between 10 and 20 to ensure balanced connectivity, and the number of frames and bins of the voxel grid was set to 16, a common configuration (Zhou et al., 2024). The results indicate that event2vec with random sampling exhibits minimal pre-processing time. While employing K-Means clustering increases pre-processing time, it still remains faster than window slicing.
Throughput and Latency Table 3 compares the throughput and single event stream inference latency of FARSE-CNN, AEGNN, SNN, Sparse GRU and our event2vec model on the DVS Gesture dataset. Our event2vec model exhibits extremely high training and inference throughput, primarily due to its full compatibility with a highly optimized linear attention library that can fully leverage GPUs for acceleration. The FARSE-CNN model exhibits the lowest throughput. This is attributable to its reliance on Sparse CNN, which receives limited acceleration from unstructured sparsity on GPUs, and the inclusion of a recurrent structure that further constrains its speed.
4.2 Ablation Experiments
Table 4: Ablation analysis of embeddings on the DVS Gesture dataset
Embedding Accuracy (%)
Spatial Temporal
[…]
Embedding Comparison We conducted an ablation study on the DVS Gesture dataset to evaluate the accuracy contributions of different components, as detailed in Table 4. We tested various combinations of spatial embedding methods (standard (Eq. 2) vs. parametric (Eq. 4)) and temporal embedding modules (sinusoidal embedding on $t$ vs. convolutional embedding on $\Delta t$). The combination of the standard embedding with our convolutional temporal embedding (Standard + Conv($\Delta t$)) yields the lowest accuracy. We attribute this to the standard embedding layer’s lack of inductive bias, which prevents it from effectively learning neighborhood semantics and subsequently hinders the convolutional temporal encoder. Consequently, when using our parametric embedding, the convolutional encoder achieves the highest accuracy. It is worth noting that our parametric embedding consistently outperforms the standard version when paired with any temporal embedding, validating the effectiveness of incorporating neighborhood semantics.
Effect of Event Numbers Processing fewer events results in lower resource consumption, which is always desirable in event-based applications. We benchmark the impact of varying the number of randomly sampled events ($L$) on several key metrics: training/inference throughput, single event stream inference latency, and accuracy on the DVS Gesture and ASL-DVS datasets. As illustrated in Figure 3, the performance trends are consistent across both datasets. For small values of $L$, throughput decreases only marginally as $L$ increases. This is attributable to the CUDA kernel launch overhead, which dominates the computation time, rendering the actual kernel execution time negligible in this regime. As $L$ grows larger, the kernel execution time becomes the primary bottleneck, causing a more pronounced decrease in throughput. Notably, while $L$ increases, the throughput decreases approximately inversely proportionally, a finding consistent with the $\mathcal{O}(L)$ complexity of linear attention. The single event stream inference latency for both models increases only slightly with $L$, further indicating that kernel launch overhead, rather than execution time, remains the dominant factor. In terms of accuracy, the general trend shows improvement as $L$ increases. Surprisingly, both models maintain a reasonable level of accuracy even with a very small number of events (e.g., $L=32$), demonstrating their robustness to sparse inputs.
Figure 3: Effect of number of events on models for (a) DVS Gesture (b) ASL-DVS.
Figure 4: DVS Gesture: Accuracy vs. number of events against Araghi et al. (2025).
We also compare our method with the sophisticated sampling techniques from Araghi et al. (2025), which use a voxel grid representation (Figure 4). The results highlight the inherent effectiveness of event2vec: when paired with simple random sampling, it consistently outperforms the voxel grid representation, even when the latter employs more complex, meticulously designed sampling strategies.
@@ 4.3 Visualization
Figure 5: Visual comparison of the learned spatial embeddings: parametric vs. standard.
@@
Event-wise Attention
@@
As event2vec is an event-wise representation, its attention mechanism can be visualized at a fine-grained, event-level resolution. Figure 6 displays attention heatmaps overlaid on the original event streams for DVS Gesture (row 1), ASL-DVS (row 2), and DVS-Lip (row 3). The visualizations reveal that the model correctly focuses on the hands in DVS Gesture, the finger joints and contours in ASL-DVS, and the lip region in DVS-Lip. However, consistent with the lower classification accuracy compared to the other two datasets, we also observe instances where the model incorrectly allocates significant attention to other facial features, such as the eyes and ears.
@@ 5 Conclusions
Neuromorphic event cameras introduce a paradigm shift in computer vision, presenting both unique opportunities and significant challenges. A central challenge has been reconciling their asynchronous, sparse nature with the synchronous, dense tensor-based architectures of deep learning. In this paper, we introduced event2vec, a novel representation that directly addresses this challenge by enabling neural networks to natively process asynchronous events. Our experimental results demonstrate that event2vec achieves accuracy competitive with established methods while offering compelling advantages in parameter efficiency, pre-processing overhead, throughput, and robustness across varying numbers of events. Beyond these performance metrics, the most significant contribution of event2vec is its conceptual alignment of event streams with the paradigm of natural language processing. This opens new avenues for research and application. By treating events as a sequential language, we can begin to explore novel applications by leveraging the sophisticated architectures developed for large language models.
@@ Acknowledgments
Reproducibility Statement
The experimental code, training logs, terminal output, and trained models for this paper are all provided together in the code repository https://github.com/Intelligent-Computing-Lab-Panda/event2vec. We have included detailed instructions in the repository, allowing readers to easily prepare the dataset and reproduce the experiments based on these instructions. All experiments in this paper fix the random number seeds of PyTorch, NumPy, and Python to 0 using the seed_everything function from PyTorch Lightning, in order to maintain reproducibility as much as possible.
@@ Appendix A Appendix
A.1 Datasets
The DVS Gesture dataset is a benchmark commonly employed for model evaluation. It comprises 11 categories of hand gestures and is officially divided into a training set with 1,176 samples and a testing set with 288 samples.
The ASL-DVS dataset contains 24 classes corresponding to letters from American Sign Language, amounting to a total of 100,800 samples. Each class consists of 4,200 samples, and the duration of each sample is approximately 100 ms. Following the methodology in Bi et al. (2019), we partition the dataset by allocating the initial 80% for training and the remaining 20% for testing.
The DVS-Lip dataset encompasses 100 word classes derived from the visual information of a speaker’s lip movements. It provides an official training/testing split with 14,896 and 4,975 samples, respectively.
A.2 Bi-directional Linear Attentions
Ignoring the layer index for simplicity, the recurrence relation for the hidden state $\textbf{S}[t]$ at time-step $t$ in a standard linear attention model is:
$\textbf{S}[t]$ $=f(\textbf{S}[t-1],\textbf{k}[t],\textbf{v}[t]),$ (8)
$\textbf{O}[t]$ $=\textbf{W}_{o}\textbf{S}[t],$ (9)
where $\textbf{k}[t]$ and $\textbf{v}[t]$ are input-dependent key and value vectors, $\textbf{W}_{o}$ is the output projection matrix, and $\textbf{O}[t]$ is the output vector. The specific recurrence function, $f$, used in this work is based on the Forgetting Transformer (FOX) (Lin et al., 2025) implemented by the Flash Linear Attention library (Yang & Zhang, 2024)
We adapt this formulation to be bi-directional by maintaining separate forward and backward hidden states, which are then concatenated and fused through a linear layer to produce the output $\textbf{O}_{fb}[t]$:
$\textbf{S}_{f}[t]$ $=f(\textbf{S}_{f}[t-1],\textbf{k}[t],\textbf{v}[t]),$ (10)
$\textbf{S}_{b}[t]$ $=f(\textbf{S}_{b}[t-1],\textbf{k}[L-t-1],\textbf{v}[L-t-1]),$ (11)
$\textbf{S}_{fb}[t]$ $=[\textbf{S}_{f}[t];\textbf{S}_{b}[t]],$ (12)
$\textbf{O}_{fb}[t]$ $=\textbf{W}_{fb}\textbf{S}_{fb}[t].$ (13)
Unlike classic bi-directional RNNs (Schuster & Paliwal, 1997) that often use independent parameters for each direction, our model employs shared parameters for the forward and backward passes. Consequently, the only increase in parameters compared to the uni-directional model arises from the output projection matrix $\textbf{W}_{fb}$, which has twice the number of parameters as the original $\textbf{W}_{o}$.
A.3 Model Hyper-parameters

### v4

@@ 3.4 Event Sampling and Aggregation
In this paper, we primarily use two methods. The first is uniform random sampling. We find that this straightforward method works well in most cases and is extremely computationally efficient. However, a significant limitation of random sampling is the substantial information loss incurred by discarding the majority of the events, leading to suboptimal accuracy in complex tasks. Our second method addresses this by leveraging K-means clustering to aggregate the entire event stream into $L$ representative clusters. Specifically, the clustering process is performed independently on the two event polarities to preserve their distinct information channels. Furthermore, we compute an intensity factor, $\rho$, equal to the number of raw events belonging to that cluster. This intensity factor then modulates the corresponding spatial embedding vector, effectively weighting the representation by its event density.
To reduce the latency of running the K-means cluster algorithm during inference, we propose a GPU-based batched K-means++ algorithm. This method approximates the step-by-step iteration of K-Means++ initialization (Arthur and Vassilvitskii, 2007) via multi-step batch computation. Meanwhile, it leverages the triangle inequality to reduce the computational cost of distance updates, and achieves an efficient GPU-based implementation with PyTorch. Details can be found in Appendix A.1.
3.5 The Formulation of Event2Vec
@@
$\textbf{V}[i]=$ $(\log\bm{\rho}[i]+1)\cdot$
$\bigg(\text{Embed}_{s}(\textbf{x}[i],\textbf{y}[i],\textbf{p}[i])+\text{Embed}_{t}(\Delta\textbf{t})[i]\bigg),$ (6)
where $\bm{\rho}$, x, y, p, and t are sequences representing the intensity factors, spatial coordinates, and timestamps of $L$ events. $\Delta\textbf{t}[i]=\textbf{t}[i+1]-\textbf{t}[i],i=\{0,1,...,L-2\}$ and $\Delta\textbf{t}[L-1]=0$. For a native event, $\bm{\rho}[i]$ is 1, while for a cluster event, it represents the number of raw events aggregated into that cluster. We take the logarithm of $\bm{\rho}$ to suppress clusters with an excessive number of events and prevent them from dominating the entire sequence.
@@ 3.6 Network Structure
The proposed methods are validated on widely used event classification tasks with a stacked network including Event2Vec, Backbone and Classification Head. Details can be found in Appendix A.2.
Event2Vec: The spatial embedding module $\phi$ consists of 3 linear layers. It gradually increases the number of features as $3\rightarrow\frac{D}{4}$, $\frac{D}{4}\rightarrow\frac{D}{2}$ and $\frac{D}{2}\rightarrow D$. Layer Normalization (Ba et al., 2016) layers are also inserted after each linear layer to stabilize training. A ReLU activation is placed after the first two Layer Normalizations. The temporal embedding module has a similar structure to the spatial embedding module, except that it replaces linear layers with depth-wise convolutional layers with a kernel size of 3 and a stride of 1, and the numbers of channels gradually increase as $1\rightarrow\frac{D}{4}$, $\frac{D}{4}\rightarrow\frac{D}{2}$ and $\frac{D}{2}\rightarrow D$.
Backbone: Backbone is constructed by stacking multiple Transformer blocks, which consists of a self-attention, a feed-forward network composed of two fully connected layers and an optional pooling to reduce the sequence length. We employ the Forgetting Transformer (Lin et al., 2025) as the self-attention in the backbone. Linear attentions such as Gated Linear Attention (Yang et al., 2024) can also be employed with slight accuracy drop, the results of which are reported in Appendix A.3. It is important to recognize that the forgetting gates in Forgetting Transformer are order-sensitive. To enhance the learning capability, we extend the Forgetting Transformer to a parameter-shared bi-directional formulation. Further details are provided in Appendix A.4.
Classification Head: We employ an average pooling layer to aggregate features across all positions in the sequence. Then a linear layer is used to make a classification decision.
@@ 4 Experiments
We conduct a series of experiments on classification tasks using three neuromorphic datasets: DVS Gesture, ASL-DVS, and DVS-Lip. In this section, results are reported in the format $a\pm b$, representing the mean and standard deviation, respectively. For experiments that involve random sampling, results are computed over 10 independent runs on the test set.
@@ 4.1 Comparison Between Representations
Accuracy and Parameter Efficiency Table 1 compares the accuracy and model parameters of event2vec with those of other representations across the three datasets. Our models for DVS Gesture and ASL-DVS are trained directly on randomly sampled events. For DVS-Lip, our model first undergoes self-supervised pre-training (refer to Appendix A.5) on cluster events. We then report the fine-tuning accuracy on both randomly sampled events and cluster events generated by the proposed Batched K-means++ algorithm. Our method achieves comparable accuracy on DVS Gesture and the highest accuracy on ASL-DVS and DVS-Lip among other leading representations, while demonstrating exceptional parameter efficiency. For example, previous State-Of-The-Art (SOTA) models use $2.79\times,815.93\times,$ and $12.22\times$ as many parameters as our model on three datasets.
Table 1: Model performance and size comparison on neuromorphic datasets.
@@
DVS Gesture Sparse GRU + Frame (Subramoney et al., 2023) 97.80 4.80
SNN + Frame (Yao et al., 2023) 98.23 6.50
@@
GNN & Transformer + Image & Voxel Graph (Yuan et al., 2023) 99.60 220.30
@@
DVS-Lip ResNet-18 & BiGRU + Frame (Tan et al., 2022) 72.1 241.20
Spiking ResNet18 & BiGRU + Frame (Dampfhoffer and Mesquida, 2024) 75.3 223.63
@@
(1024 Batched K-Means++ Cluster Events) 75.88
Table 2: Throughput and latency comparison between Event2Vec and previous SOTA models.
Dataset Method Batch Throughput (samples/s) Single-stream Inference
Training Inference Latency (ms) GPU Memory (MB)
Data Pre-processing Forward Total
DVS Gesture Max-Former 241.12 $\pm$   0.55 1077.35 $\pm$       2.20 10.29 $\pm$   0.23 23.89 $\pm$ 11.65 34.18 $\pm$ 11.75 834
Event2Vec 1016.19 $\pm$ 61.18 2900.08 $\pm$   277.30 9.92 $\pm$   5.79 13.51 $\pm$   9.04 23.43 $\pm$ 10.63 602
ASL-DVS GNN & Transformer 78.08 $\pm$   6.39 200.16 $\pm$     12.66 55.12 $\pm$ 38.35 10.90 $\pm$   0.45 66.02 $\pm$ 38.41 5464
Event2Vec 933.58 $\pm$ 16.14 12543.81 $\pm$ 2565.54 1.10 $\pm$   0.23 6.24 $\pm$   2.89 7.34 $\pm$   2.91 824
DVS-Lip Spiking ResNet18 & BiGRU 10.85 $\pm$   0.03 165.29 $\pm$       2.25 373.82 $\pm$   2.21 15.07 $\pm$   7.73 388.89 $\pm$   8.14 1226
Event2Vec 383.71 $\pm$   0.89 942.29 $\pm$     22.36 17.23 $\pm$   3.33 39.87 $\pm$   1.28 57.10 $\pm$   3.62 838
Throughput, Latency and Memory We further compare model throughput and latency between our models and previous SOTA models, and results are shown in Table 2. Appendix A.6 provides more details about these experiments. Throughput is a primary metric governing the efficiency of model training. For different models, we set the batch size to the largest possible power of 2 or the average of two adjacent powers of 2 (e.g., 64, 96, 128, …) without exceeding the GPU memory limit, aiming to maximize the reported throughput. The results demonstrate that event2vec fully leverages the computational efficiency of Transformers, achieving training and inference throughput that is $4.21\times$ and $2.69\times$, $11.96\times$ and $62.67\times$, and $35.36\times$ and $5.70\times$ higher than those of prior works across the three datasets, respectively. For inference tasks on edge devices (e.g., an embedded neuromorphic system), the latency of processing a single event stream and the GPU memory consumed by the model are critical. We conducted comparative experiments and measured three latency components: event data pre-processing latency, model forward propagation latency, and total latency. The results show that across the three datasets, the total latency of our model is only $68\%,11.12\%$, and $14.68\%$ of that of the prior SOTA methods, while the memory consumption is merely $72.18\%,15.08\%$.
4.2 Ablation Experiments
Embedding Comparison We conducted an ablation study on the DVS Gesture dataset to evaluate the accuracy contributions of different components, as detailed in Table 4. We tested various combinations of spatial embedding methods (standard (Eq. 2) vs. parametric (Eq. 4)) and temporal embedding modules (sinusoidal embedding on t vs. convolutional embedding on $\Delta\textbf{t}$). The combination of the standard embedding with our convolutional temporal embedding (Standard + Conv($\Delta\textbf{t}$)) yields the lowest accuracy. We attribute this to the standard embedding layer’s lack of inductive bias, which prevents it from effectively learning neighborhood semantics and subsequently limits the performance of the convolutional temporal encoder. Consequently, when using our parametric embedding, the convolutional encoder achieves the highest accuracy. It is worth noting that our parametric embedding consistently outperforms the standard version when paired with any temporal embedding, validating the effectiveness of incorporating neighborhood semantics.
Robustness to the Number of Events Processing fewer events results in lower resource consumption, which is always desirable in event-based applications. Figure 2 compares our method with the sophisticated sampling techniques from (Araghi et al., 2025), which use a voxel grid representation. The results highlight the inherent effectiveness of event2vec: when paired with simple random sampling, it consistently outperforms the voxel grid representation, even when the latter employs more complex, meticulously designed sampling strategies. Appendix A.7 provides additional details on the variations in the performance of the event2vec model with respect to the number of events.
Figure 2: Accuracy vs. number of events compared to sampling techniques from (Araghi et al., 2025) on the DVS Gesture dataset.
Figure 3: Accuracy vs. spatial resolution: Comparison with the SOTA Max-Former (Fang et al., 2025) on DVS Gesture.
Robustness to Spatial Resolutions We further tested the robustness of event2vec to variations in the spatial resolution of sensors and compared it with Max-Former (Fang et al., 2025) on the DVS Gesture dataset. For event2vec, the coordinates are treated as floating-point numbers, scaled to the target resolution $h\times w$ and then quantized. Subsequently, the coordinates are upscaled back to the original resolution $H\times W$ and re-quantized. For Max-Former, which adopts a frame-based representation, bilinear interpolation is used to downscale the frames to a lower resolution $h\times w$, followed by upscaling back to the original resolution $H\times W$. The results in Figure 3 demonstrate that event2vec exhibits strong robustness to changes in resolution. Moreover, it maintains a classification accuracy significantly higher than random guessing even when the resolution is reduced to $1\times 1$ (i.e., complete loss of spatial information), which indirectly validates the effectiveness of the temporal embedding.
[…]
Table 3: Ablation analysis of embeddings on DVS Gesture.
Spatial Embedding Temporal Embedding Accuracy (%)
@@ 4.2 Ablation Experiments
Table 4: K-Means clustering latency and accuracy on DVS-Lip.
Method Latency (ms) Accuracy (%)
[…]
Scikit-learn (iters=300) 383.42 $\pm$ 149.22 75.08
Scikit-learn (iters=20) 374.22 $\pm$ 145.97 74.72
@@
Cluster Latency Table 4 compares the average per-sample clustering latency on the test set, as well as the test-set accuracy of models trained with clustered data, when different K-Means clustering methods are applied to the DVS-Lip dataset. We compare our approach with two benchmarks: Scikit-learn’s CPU-based K-Means (using K-Means++) and Meta’s Faiss GPU K-Means (Johnson et al., 2019).As shown in Table 4, the proposed batched K-Means++ outperforms benchmarks by achieving peak accuracy with minimal latency. Notably, it is significantly faster than Scikit-learn and more accurate than the GPU-accelerated Faiss. While Faiss (iters=20) offers comparable speed, it suffers a $1.48\%$ accuracy drop compared to our approach.
@@ 4.3 Visualization
Figure 4: Visual comparison of the learned spatial embeddings.
@@
Event-wise Attention As event2vec is an event-wise representation, its attention mechanism can be visualized at a fine-grained, event-level resolution. Figure 5 displays attention heatmaps overlaid on the original event streams for DVS Gesture (row 1), ASL-DVS (row 2), and DVS-Lip (row 3). The visualizations reveal that the model correctly focuses on the hands in DVS Gesture, the finger joints and contours in ASL-DVS, and the lip region in DVS-Lip. However, consistent with the lower classification accuracy compared to the other two datasets, we also observe instances where the model incorrectly allocates significant attention to other facial features, such as the eyes and ears.
@@ 5 Conclusions
Neuromorphic event cameras introduce a paradigm shift in computer vision, presenting both unique opportunities and significant challenges. A central challenge has been reconciling their asynchronous, sparse nature with the synchronous, dense regular architectures of deep learning. In this paper, we introduced event2vec, a novel representation that directly addresses this challenge by enabling neural networks to natively process asynchronous events. Our experimental results demonstrate that event2vec achieves accuracy competitive with established methods while offering compelling advantages in parameter efficiency, pre-processing overhead, throughput, and robustness across varying numbers of events and spatial resolutions. The remarkable efficiency and robustness of event2vec suggest its significant potential for real-time deployment on resource-constrained edge devices, where low-latency sensing and low-power consumption are paramount. Beyond these performance metrics, the most significant contribution of event2vec is its conceptual alignment of event streams with the paradigm of natural language processing. This opens new avenues for research and application. By treating events as a sequential language, we can begin to explore novel applications by leveraging the sophisticated architectures developed for large language models.
@@ Acknowledgments
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Appendix
A.1 Batched K-Means++ Cluster Algorithm
For challenging classification tasks like DVS-Lip, random sampling leads to significant information loss, resulting in suboptimal accuracy. To ensure that all events contribute to the final representation, we employ an event clustering approach. Given a raw event stream $\mathcal{E}=\{(\mathbf{x},\mathbf{y},\mathbf{t},\mathbf{p})\}$ containing $N$ events, our objective is to generate $L$ cluster event streams $\mathcal{R}=\{(\mathbf{x}_{c},\mathbf{y}_{c},\mathbf{t}_{c},\mathbf{p}_{c},\bm{\rho})\}$, where $(\mathbf{x}_{c},\mathbf{y}_{c},\mathbf{t}_{c},\mathbf{p}_{c})$ denotes the coordinates of the cluster centers and $\bm{\rho}$ represents the event count within each cluster. Notably, we perform clustering separately for the two polarities to prevent the loss of physical significance caused by polarity mixing. Specifically, assuming the number of events for each polarity is $N_{0}$ and $N_{1}$, we calculate the numbers of clusters as $L_{0}=\text{Round}(\frac{N_{0}}{N}\cdot L)$ and $L_{1}=L-L_{0}$, respectively. Finally, the two resulting cluster streams are merged and sorted chronologically.
The general practice of K-Means clustering involves using the K-Means function provided in Scikit-learn (sklearn)(Pedregosa et al., 2011), a Python machine learning library. However, this function is implemented on the CPU, resulting in slow execution when the number of events is large, which significantly increases the latency of the model in processing real-time tasks. To address this issue, we propose a GPU-based Batched K-Means++ Event Clustering algorithm, whose detailed workflow is presented in Algorithm 1. The acceleration of this algorithm mainly stems from the following optimizations:
  1. 1.
    Reduced Loop Overhead: The number of loop iterations is reduced by a factor of $B$.
  2. 2.
    Parallel Computing: Utilizes torch.cdist to compute distances from all points to the batch of $B$ new centers in parallel.
  3. 3.
    Incremental Update: The update of $\mathbf{D}^{2}$ leverages a variant of the triangle inequality, requiring comparison only between the current known minimum distance and the distance to the newly added batch of centers, avoiding recomputation of all pairwise distances.
1
Input: Raw Event Stream $\mathcal{E}=\{(\mathbf{x},\mathbf{y},\mathbf{t},\mathbf{p})\}$, Total Points NN
Params: Target Cluster Count $L$, Spatial Dimensions $H,W$, Batch Size $B$, Max Iterations $I_{max}$
Output: Downsampled Event Set $\mathcal{S}$
2
/* 1. Data Preprocessing & Normalization */
3 Move the data to the GPU
4 Compute normalized time $\hat{\mathbf{t}}=\frac{\mathbf{t}-t[0]}{t[N-1]-t[0]}$
5 Construct 3D feature space point set $\mathbf{V}=\{(\frac{x[i]}{W},\frac{y[i]}{H},\hat{t}[i])\}_{i=0}^{N-1}$
6 Split $\mathbf{V}$ into positive set $\mathbf{V}^{0}$ and negative set $\mathbf{V}^{1}$ based on polarity $\mathbf{p}$
7 Allocate target center counts $L_{0}$ and $L_{1}$
8 Initialize result set $\mathcal{R}=\emptyset$
9
/* Cluster for each polarity separately */
10 for each point set $\mathbf{V}_{sub}\in\{\mathbf{V}^{0},\mathbf{V}^{1}\}$ and target count $K_{sub}\in\{L_{0},L_{1}\}$ do
11 if $|\mathbf{V}_{sub}|==0$ then
12    continue
13   end if
14
 /* Phase 1: Batched K-Means++ Initialization */
15   Randomly select the first center $\mathbf{c}_{0}\in\mathbf{V}_{sub}$, initialize center set $\mathbf{C}=\{\mathbf{c}_{0}\}$
16   Compute squared distance from all points to first center $\mathbf{D}^{2}=\|\mathbf{V}_{sub}-\mathbf{c}_{0}\|^{2}$
17
18 while $|\mathbf{C}|<K_{sub}$ do
19      Calculate sample count for this batch $M=\min(B,K_{sub}-|\mathbf{C}|)$
    /* Parallel Sampling: Use current distance as probability weights */
20      Sample $M$ new candidate centers $\mathbf{C}_{new}$ based on weights $\mathbf{w}\propto\mathbf{D}^{2}$
21      Add $\mathbf{C}_{new}$ to $\mathbf{C}$
    /* Incremental Distance Update: Only to newly added centers */
22    $\mathbf{D}^{2}_{new}=\min_{c\in\mathbf{C}_{new}}\|\mathbf{V}_{sub}-c\|^{2}$
23      Update global minimum distance $\mathbf{D}^{2}\leftarrow\min(\mathbf{D}^{2},\mathbf{D}^{2}_{new})$
24
25   end while
26
 /* Phase 2: Standard Lloyd’s Iteration */
27 for $iter=0$ to $I_{max}-1$ do
28    E-step: Assign labels to points based on the nearest center in $\mathbf{C}$
29    M-step: Compute centroids of each cluster as the new $\mathbf{C}$
30    if the center shift <tol<tol then
31       break
32      end if
33
34   end for
35
 /* Denormalization & Intensity Calculation */
36   Count points in each cluster as Intensity $\bm{\rho}$
37   Map coordinates of $\mathbf{C}$ back to physical dimensions $(W,H,t_{span})$
38   Add result $(\mathbf{x}_{c},\mathbf{y}_{c},\mathbf{t}_{c},\mathbf{p}_{sub},\bm{\rho})$ to $\mathcal{R}$
39
40 end for
41
/* 3. Post-processing */
42 Merge all results in $\mathcal{R}$
43 Sort results by time $\mathbf{t}_{c}$ to ensure memory contiguity
44 Return Sorted Tensor Data
Algorithm 1 Batched K-Means++ Event Clustering on GPU
A.2 Model Structures and Hyper-parameters